ill save you S3,000 ın he first year. The machine will at the savings decline a a rate of 2% per ha s he present value of h You are thinking o building a new machine ha savings if the interest rate is 5% per year? hen begn to wear out so ar o ever The present value of the savings is SRound to the nearest dollar.)
Answer #1

Value after first year=(Savings in first year(1+Growth rate))/(Interest Rate-Growth Rate)

=(3000(1-0.02))/(0.05-(0.02))

=(2940/0.07)

=$42000

Hence present value of savings=Future value*Present value of discounting factor(rate%,time period)

=3000/1.05+$42000/1.05

which is equal to

=$42857(Approx).